Presentation is loading. Please wait.

Presentation is loading. Please wait.

CALICE WHCAL testbeam at SPS H8 27 sep – 03 oct: 6 days for energies up to 180 GeV (+ polarity)

Similar presentations


Presentation on theme: "CALICE WHCAL testbeam at SPS H8 27 sep – 03 oct: 6 days for energies up to 180 GeV (+ polarity)"— Presentation transcript:

1 CALICE WHCAL testbeam at SPS H8 27 sep – 03 oct: 6 days for energies up to 180 GeV (+ polarity)

2 Setup 10 oct LCD monthly Erik van der Kraaij, CERN LCD2 TCMT: Steel/scintillator-strip sandwich calorimeter with 16 layers (~5.5 λ) Current HCAL setup has 38 W layers. Including active material this is ~4.8 λ

3 Repairs before the run Week before the run, Ivo Polak came to help tune the Calibration and Monitoring Board #20 –Six boards however have to be taken out to reach #20. –After tuning we retested all  all OK. Weekend before the run, Kurt Francis helped with the tailcatcher –MIPs for TCMT strips –Gains –Mapping problem 10 oct LCD monthly Erik van der Kraaij, CERN LCD3

4 TCMT mapping 10 oct LCD monthly Erik van der Kraaij, CERN LCD4

5 TCMT mapping 10 oct LCD monthly Erik van der Kraaij, CERN LCD5 Best guess why the map is wrong: there are 2 different boxes for cabling

6 TCMT mapping - fixed 10 oct LCD monthly Erik van der Kraaij, CERN LCD6

7 Muon samples 10 oct LCD monthly Erik van der Kraaij, CERN LCD7 For MIP calibration and TCMT mapping: In the days before & after our beam time, we scanned the full surface in 9 positions. > 200k muons in each position.

8 High energy hadron program Got help from Ilias & Simon on better beam files Took 200k events at: 180, 150, 120, 100 GeV (secondary beam) Took 400k events at: 180, 150, 120, 100 GeV (tertiary beam) Took 4M events at 60 and 4M at 80 GeV:  ~ Kaon events at each energy. 10 oct LCD monthly Erik van der Kraaij, CERN LCD 8 Secondary beam runs September vs July

9 Low energy program Friday 6 pm - Saturday 8 am: got wobbling set to 50 GeV (with 80 GeV for ATLAS in H6) Pion samples (200k each) at 30 & 50 GeV Electron samples (200k each) at 15, 20 and 30 GeV 10 oct LCD monthly Erik van der Kraaij, CERN LCD 9 For the electron runs at 15 & 20 GeV, took sample with & without all scintillators along the beamline  Perhaps an explanation for the long tails we see in June/July samples. With scint. Without scint.

10 Problems during run Saturday morning TCMT 1B started drawing high current on the SiPM. Replaced total FE board  Did not help. Other monitoring tools looked fine  did not touch it anymore. Forgot to restart LEDs at this point 3 gain runs (2 days) are empty. 10 oct LCD monthly Erik van der Kraaij, CERN LCD 10 For three runs layer #33 suddenly gave too many hits and too high energy response. Do not know why… Restarted these runs. Not related to LEDs being off, it also happened a day before TCMT problem.

11 Conclusion 10 oct LCD monthly Erik van der Kraaij, CERN LCD11 Beam operation went well, got help for better focussing. Detector repairs before run were successful Some malfunctioning during run: –TCMT 1B –Layer #33 We completed full 2011 program. –11 M events in pion runs. –2.2 M events in dedicated muon runs. –1.1 M events in electron runs.

12 Temperatures From the T3B team we got a device with 21 Pt_1000 temperature sensors, easily read out via USB. Placed these on a 1 m 2 aluminium plate  slided in the free slot #39; Goal is to obtain temperature profile over the surface, vs time. 10 oct LCD monthly Erik van der Kraaij, CERN LCD12 We see the day/night fluctuations, but variations among sensors is max 0.6 °C. The free slot might be too far from electronics  We’re repeating this measurement now in slot #20.

13 Backup

14 TCMT Front End board 39B Saturday morning TCMT 1B started drawing high current on the SiPM. Unplugged all cables Replaced several FE chips Nothing worked. After this, could also not configure chips anymore.  Replaced total FE board 10 oct LCD monthly Erik van der Kraaij, CERN LCD 14 Could reconfigure again, high current remains. All other monitoring tools looked fine.  did not touch it anymore. Unfortunately forgot to reset LEDs for AHCAL! The 4 gain runs from this point on are empty… (2 days) μA

15 HCAL layer #33 For three runs layer #33 suddenly gave too many hits and too high energy response. Do not know why… Restarted these runs. Not related to LEDs being off, it also happened a day before TCMT problem. 10 oct LCD monthly Erik van der Kraaij, CERN LCD 15 μA

16 CMB #20 fine tuning Week before the run, Ivo Polak came to help tune the Calibration and Monitoring Board #20 –Via LEDs a CMB sends light to the tiles for calibration. –#20 never functioned properly since installation at CERN –Each of the 11 LEDs has a potentiometer to be tuned.  LED #8 could not be tuned, light intensity fluctuated Six boards however have to be taken out to reach #20. After tuning we retested all six –#20 now finally sees light: for ~70% of channels the gain can be extracted –No significant effect on other 5 modules. 10 oct LCD monthly Erik van der Kraaij, CERN LCD16

17 TCMT MIP peaks The weekend before the run, Kurt Francis visited CERN and helped with the tailcatcher. Extended AngleTrackFinder to find mips in TCMT –extrapolate track to TCMT and allow +- 1 strip Additional quality cuts –Consider only hits which form a tower of at least 5 in the TCMT alone  Some channels give nice results, however still have the MIP peak very close to the pedestal 10 oct LCD monthly Erik van der Kraaij, CERN LCD17

18 First night – secondary beam 200k events at 180, 150, 120 and 100 GeV. Compared to same energies (but negative polarity) runs of July  Linearity is good Higher energies of new run are probably due to lower temperatures 10 oct LCD monthly Erik van der Kraaij, CERN LCD 18

19 TCMT MIP peaks 10 oct LCD monthly Erik van der Kraaij, CERN LCD19


Download ppt "CALICE WHCAL testbeam at SPS H8 27 sep – 03 oct: 6 days for energies up to 180 GeV (+ polarity)"

Similar presentations


Ads by Google